OpenStudy (anonymous):

Write a mathematical expression for the statement 'Y varies jointly with x and the cube of z, and inversely with the square root of w' Could someone help me understand what this one wants? I have no clue what the 'varie' means

OpenStudy (anonymous):

the term is "vary" when something varies with another, it means they change depending on the other if x varies inversely to y, then \[x = \frac{ 1 }{ y }\]

OpenStudy (anonymous):

Do you need more help?

OpenStudy (anonymous):

if x varies jointly with y, then x is directly related to y

OpenStudy (anonymous):

I think I'm sort of getting it. Then X would = Y, since they're directly related?

OpenStudy (anonymous):

Yup So what would the overall equation be?

OpenStudy (anonymous):

Hm, we would write it as one equation? I would think it would be three seperate ones since Y changes. Y=X \[Y=z^{2}\] then\[Y^{-1}= \sqrt{w}\]

OpenStudy (anonymous):

Ah but the original equation said that 'Y varies jointly with x and the cube of z, and inversely with the square root of w' Note "and" This means that y varies with x,z^2 and rt(w) in one equation

OpenStudy (anonymous):

I don't know how better to explain that, I was only using y=x to demonstrate y varies jointly with x

OpenStudy (anonymous):

Hm,ok. I'll see how I can work that That helps so much! Not only with this question but with my whole paper. Thank you!

OpenStudy (anonymous):

So, whats the equation?

OpenStudy (anonymous):

For some questions that require you to input numerical data, you will need to include a "constant" \[y = \frac{ kxz ^{2} }{ \sqrt{w} }\]

OpenStudy (anonymous):

k = constant

OpenStudy (anonymous):

Oh, I see! The square of W goes on the bottome because it's the inverse. and the 'and' simply means they are multiplied together. For some reason I thought they would be added instead.